# Internal Consistency and Self-Feedback in Large Language Models: A Survey

### v2

@@ Abstract
In this paper, we summarize a theoretical framework, Internal Consistency, offering explanations for reasoning deficiencies and hallucinations. Internal Consistency refers to the consistency in expressions among LLMs’ latent, decoding, or response layers based on sampling methodologies. Then, we introduce another effective theoretical framework capable of mining Internal Consistency, named Self-Feedback. This framework consists of two modules: Self-Evaluation and Self-Update. The former captures Internal Consistency Signals, while the latter leverages the signals to enhance either the model’s response or the model itself. This framework has been employed in numerous studies.
@@ I Introduction
Large language models (LLMs) have made significant advancements in the field of natural language processing (NLP), demonstrating capabilities close to basic human intelligence. These include features like performing basic reasoning and learning from examples [1].
However, LLMs still face many challenges. For instance, current models struggle to generate consistent responses [2], exhibit illogical reasoning when faced with out-of-distribution problems [3], and display excessive confidence with insufficient awareness of their own capability boundaries [4].
Among the numerous issues, we identify a fundamental category, Internal Consistency, as being central to the core challenges. At the surface level, even robust language models such as GPT-4o generate inconsistent responses to identical queries. This phenomenon is illustrated using GPT-4o as an example, as depicted in Fig. 1. At the middle layer, the inconsistency is caused by using trivial stochastic sampling methods (Top-k, Top-p, beam search, etc.) during decoding. At the deepest layer,  [5, 6, 7] point out that specific attention heads in latent states are related to the faithfulness of the responses, and increasing their weights can consistently enhance the faithfulness of the answers.
Figure 1: GPT-4o provides different answers to the same question. The complete responses can be found in our GitHub open-source repository.
“Model consistency is weak” is closely related to the topics such as “models lack reasoning” and “models exhibit hallucinations.” We believe that Internal Consistency is a more profound topic underlying issues of reasoning and hallucinations. Simply put, only when the model exhibits internal inconsistency can we employ strategies like Self-Consistency [2], Self-Refine [8], and Self-Correct [9] to elevate LLMs’ reasoning abilities and alleviate the hallucinations. We refer to these strategies as “Internal Consistency Mining.”
@@
Internal Consistency Mining refers to developing methods to ensure Large Language Models consistently express their understanding learned from corpus.
Methods mentioned in the definition include response strategies (e.g., Chain-of-Thought, or CoT [10]), decoding strategies (e.g., Self-Evaluation Decoding [11]), and latent state strategies (e.g., Inference-Time Intervention [5]).
@@ I-A Lack Reasoning and Exhibit Hallucination
The issues of “lack reasoning” and “exhibit hallucinations” in models represent persistent concerns, with their prominence in the academic community demonstrably increasing, as illustrated by Google Trends [^1] https://trends.google.com/ data shown in Fig. 2. In this section, we compare these two pivotal issues to demonstrate the necessity of examining them from the perspective of Internal Consistency. The definitions of two issues, real-world examples222 [12] provides more similar examples, which are simple for humans but challenging for LLMs. This work also indicates that the emergent reasoning abilities of the models might only be rudimentary and mnemonic., and benchmark examples are shown in Table I.
Figure 2: Relative search interest for the keywords “LLM Hallucination” and “LLM Reasoning.” The data retrieval date is June 14, 2024.
TABLE I: Definitions, Real-World Example, and Benchmark Examples.
@@
Real-World Example User: There is a family where mom and dad have 3 daughters and each daughter has 1 younger brother, so how many people are there in this family? GLM-3: This family consists of: (1) Mom and Dad: 2 people; (2) 3 daughters: 3 people; (3) Each daughter has 1 younger brother, and since there are 3 daughters, this means there are 3 younger brothers: 3 people. Adding these numbers together, there are a total of 2 (mom and dad) + 3 (daughters) + 3 (younger brothers) = 8 people in this family.
@@
“Reasoning” and “hallucination” have different connotations; however, it is challenging to clearly distinguish between “lack reasoning” and “exhibit hallucinations.” As illustrated in Table I, in the real-world example, an LLM assumes that each daughter has a corresponding brother, clearly demonstrating a misunderstanding of the question—a hallucination that contradicts the user’s input. This can also be interpreted as a failure in reasoning, as the model doesn’t deduce that “each having a brother means there is only one brother in total.” Consequently, it is difficult to definitively determine whether this scenario is a hallucination or a lack of reasoning. Similarly, MMLU [15] serves as a widely recognized reasoning evaluation benchmark, while TruthfulQA [16] is a hallucination evaluation benchmark. Yet, both benchmark examples in Table I, addressing financial topics in a question–answering format, make it even harder to find an essential difference between them.
Consequently, we argue that “lack reasoning” and “exhibit hallucinations” share the same essence. However, due to the different keywords, many works claim their methods solely “elevate reasoning”, even though they compare against baselines claiming to “alleviate hallucinations”, and vice versa. For example, Zhang et al. [17] proposed a method named RATT to enhance the reasoning ability but used the hallucination evaluation benchmark TruthfulQA [16] in the experiments. Similarly, Zhang et al. [7] focused on exploring methods to reduce hallucinations but employed the TriviaQA benchmark [18], supposedly to “test the model’s reasoning ability”. These instances underscore the confusion in the development of this field caused by the misuse of terminology.
Therefore, a unified perspective is required to analyze these two similar phenomena. We adopt “Internal Consistency Mining” as a term to encompass methods aimed at “reasoning elevation” and “hallucination alleviation”. Indeed, LLMs don’t comprehend reasoning or hallucinations; they only predict the next token based on probabilistic principles.
@@ I-B Self-Feedback to Promote Internal Consistency
To enhance the model’s internal consistency, scaling the model’s parameters is the most direct approach [19]. However, this doesn’t address the fundamental problem of weak consistency, as shown in Fig. 1. Furthermore, small language models provide distinct advantages, including lower computational costs and edge-side availability [20]. This indicates that while studying scaling models, it is also crucial to explore strategies for maximizing the capabilities given a small language model.
Numerous initiatives have been undertaken to improve the Internal Consistency of models. A pivotal approach involves mimicking human thought processes, which enables models to self-evaluate their own outputs and self-update their structure or outputs. Notable examples include Self-Consistency [2], which prompts the model to generate multiple answers to check for consistency (Self-Evaluation), and then use a majority voting strategy to select the final answer (Self-Update), thereby enhancing reasoning capabilities. Another example is Self-Contradict [21], which induces models to generate diverse content and checks for contradictions (Self-Evaluation), allowing the model to resolve contradictions autonomously (Self-Update) to reduce hallucinations.
This paradigm exhibit high scalability. During Self-Evaluation, it is possible to not only inspect the model’s responses but also examine its logits and the latent states. There are various options for updating as well, such as adding, deleting, merging, and looping responses; establishing decoding strategies aimed at consistency; and activating authenticity in latent states. We refer to the Self-Evaluation and Self-Update paradigm as Self-Feedback.
@@
I-C Related Surveys
Several other surveys have also focused extensively on works prefixed with “Self-”. To our knowledge, surveys such as [22],  [23], and  [24] demonstrate substantial similarities to our work. We present a straightforward comparison in Table II.
A Survey on Self-Evolution of Large Language Models [22] primarily focuses on two aspects of literature. It covers articles on LLMs generating their own training data and papers employing multi-agent approaches for iterative optimization. Compared to other studies, this study is the most comprehensive in content. As shown in Table II, the methods summarized in this survey encompass various tasks, including Instruction Following, Code Generation, and Planning. Given the wide range of tasks covered, the objectives of Self-Evolution proposed in this survey may lack a clear focus.
Automatically Correcting Large Language Models: Surveying the Landscape of Diverse Automated Correction Strategies [23] specifically concentrates on Self-Correction, a process where models correct their own errors. Defined tasks encompass reasoning errors and harmful information. Owing to its explicit task definition, this survey offers a more detailed and comprehensive theoretical analysis. These tasks are classified based on the timing of correction, including training-time, generation-time, and post-hoc corrections. Although the classification approach is logical, it doesn’t provide a comprehensive overview related to latent layers. Furthermore, the survey introduces a third task, Biased and Harmful Content Elimination, which is notably different from the first two tasks and leans towards subjective evaluation, indicating the potential for improvement in the task definition.
When Can LLMs Actually Correct Their Own Mistakes? A Critical Survey of Self-Correction of LLMs [24] also focuses on Self-Correction, delving deeper into the issue by primarily questioning whether models can genuinely Self-Correct. However, the survey’s approach to this question is limited as it only considers cases where Feedback Signals are textual and partially external, thereby directly constraining the scope of Self-Correction. Consequently, we find that the survey’s summary is not sufficiently comprehensive and lacks adequate support for its conclusions. In Section IX, we further explore this issue and provide a more insightful analysis.
@@ I-C Related Surveys
    Internal Consistency theoretical framework. We offer an in-depth review of LLMs’ Internal Consistency, examining its phenomena, formalization, status quo, etc. Furthermore, we introduce the task of Internal Consistency Mining, providing a unified perspective for reasoning elevation and hallucination alleviation tasks.
@@
    Self-Feedback theoretical framework. Our framework includes Self-Evaluation, Consistency Signal Acquisition, and Self-Update. Characterized by its simplicity and comprehensiveness, this framework is poised to inspire further research. We summarize a broad array of Self-Evaluation strategies that extend from model responses to latent space exploration. These strategies allow us to capture a diverse range of Feedback Signals, extending beyond the scalar, textual, and external signals discussed in other surveys, to include contrastive signals.
@@
Additionally, we also incorporates insights from other weakly related surveys. Section IV-A draws inspiration from [25], which meticulously investigates the uncertainty issues in LLMs. Additionally, [26] offers in-depth analyses of strategies like Chain of Thought, Self-Consistency, Tree of Thought, and Graph of Thought, which have guided our discussions in Section V-A. Moreover, the methodologies and insights from surveys on knowledge distillation [27] and preference learning [28] have been valuable in shaping Section VII.
I-D Structure of Our Work
The logical structure and the organizational framework of this paper are depicted in Fig. 3. Our research begins with the existing problem of low Internal Consistency in LLMs (Section II-C). Specific manifestations of low Internal Consistency include poor reasoning capabilities in question-answering (QA) scenarios and hallucinations in free-form generation (Section I-A). From a causal perspective, elements contributing to low Internal Consistency include inadequate latent reasoning abilities, the snowball effect of hallucinations, and the stochastic parrot hypothesis (Section II-D). We formalize internal consistency as the sampling-based consistency of model expressions across different layers (Section II-A). This involves enhancing response consistency, decoding consistency, and latent consistency (Sections II-AII-B).
To improve Internal Consistency, we propose Internal Consistency Mining across these layers. While scaling up the model is an intuitive solution, it comes with various cost-related challenges. Thus, we focus on the Self-Feedback theoretical framework, which mainly includes Self-Evaluation, Consistency Signal Acquisition, and Self-Update. Models obtain different forms of Internal Consistency signals through Self-Evaluation, and subsequently use these signals to Self-Update either responses or the model itself (Section III). We explore six lines of work in Consistency Signal Acquisition (Section IV) and seven lines of work utilizing the Self-Feedback framework, divided into three lines dedicated to reasoning elevation (Section V) and four lines aimed at hallucination alleviation (Section VI).
@@ I-E Out-of-scope Topics
  • While this paper focuses on reasoning elevation and hallucination alleviation, we don’t systematically review these aspects. Instead, we primarily discuss articles that utilize the Self-Feedback framework.
  • The methods discussed in this paper aim to enhance the Internal Consistency. Although some methods may also employ the Self-Feedback framework, if their primary goal is not to improve Internal Consistency, they remain outside the scope. For example, DINO [29] focuses on generating datasets to train better Embedding models.
  • The survey exclusively explores internal consistency and doesn’t examine the interplay between internal and external consistencies. For instance, if a user maintains an opinion contrary to the LLM, should the LLM maintain its high internal consistency? This raises questions about the autonomy of models, which are profound but beyond our current scope. We proceed assuming agreement on correctness between the model and the user.
  • Consistent with many related surveys summarizing methods whose names beginning with “Self-”, our focus is on the model’s self-awareness, self-assessment, self-correction, etc. The methods reviewed advocate for a model-in-the-loop approach, with minimal human intervention during the Self-Evaluation and Self-Update.
  • Although retrieval-augmented generation (RAG) is notable for mitigating external hallucinations [30], this paper doesn’t actively discuss RAG and primarily considers hallucinations caused by internal consistency to explore the limits of model honesty.
@@ II-A Formulation
Consistency is a critical term in logic, referring to a system where no two statements contradict each other [31]. However, systems like those of language models typically exhibit inconsistencies. For instance, as shown in Fig. 1, even GPT-4o can’t guarantee consistent responses to the same question. To better define the internal consistency, we utilize a sampling-based approach to model expressions in LLMs [25]. In addition, Table III provides explanations of some notations frequently used in this paper.
@@
[…]
@@
  • Expression from Latent Layer (vector). Expression at this layer encompasses the different activation of attention heads and latent states across the model’s architecture, contributing to diverse outputs.
For the expression type $e$, the distribution of different expressions produced by model $\mathcal{M}$ in response to the query $\bm{x}$ can be defined as follows:
@@
By sampling from this distribution (e.g., repeatedly prompting the model with a query or its synonymous variants), we can obtain a sampling set with potentially repeated elements:
@@
We will comprehensively discuss existing methods for acquiring consistency signals in Section IV.
@@ II-B The Hourglass Evolution of Internal Consistency
Response Layer. We used common Top-p sampling decoding with a temperature coefficient control to sample five times. To induce the model to output longer and more diverse content, we enabled CoT reasoning. We observed the model’s final text choices during free generation. One example output is: “Let’s think step by step. There is one period at the end of the first part, then another after the second part, and finally one more after the third part. So, there are 3 periods in total.” The resulting simplified sampling set is $\mathcal{Y}_{\text{response}}=\{5,3,3,3,3\}$.
@@
The experimental results are also shown in Fig. 5. We observed that from the shallow layers of the latent layer to the deeper layers, through the intermediate decoding layer, and finally to the response layer, the consistency of the model’s answers exhibits an “hourglass evolution” pattern.
@@
We analyze this phenomenon as follows: In the latent state, since the forward propagation is not yet complete, the attention heads near the bottom layers may tend to choose answers randomly. In contrast, the attention heads near the top layers can continually accumulate knowledge from the previous layers due to the presence of residual connections, leading to a gradual convergence in judgment and increased certainty in answers. During the decoding phase, all decoding strategies tend to output the token with the highest probability, thus maintaining high certainty. However, at the response stage, greater variability appears. We believe this is because, when the LLM generates the first token, it has already conducted reasoning (latent reasoning [^6] Yang et al. [33] introduce latent reasoning. This concept implies that when a Transformer model responds to a query, it has already conducted reasoning before generating the first token, similar to how humans think before speaking. We will also explain this concept further in Section II-D.) and make an initial judgment of the answer. However, during the response phase, the output tokens such as “I’m willing to help.” can interfere with the model’s initial reasoning and preliminary judgment, leading to a collapse in the computation results during latent reasoning.

### v3

@@ Abstract
In this paper, we use a unified perspective of internal consistency, offering explanations for reasoning deficiencies and hallucinations. Internal consistency refers to the consistency in expressions among LLMs’ latent, decoding, or response layers based on sampling methodologies. Then, we introduce an effective theoretical framework capable of mining internal consistency, named Self-Feedback. This framework consists of two modules: Self-Evaluation and Self-Update. The former captures internal consistency signals, while the latter leverages the signals to enhance either the model’s response or the model itself. This framework has been employed in numerous studies.
@@ I Introduction
Large language models (LLMs) have significantly advanced natural language processing (NLP), showing near-human capabilities in reasoning and learning from examples [1]. However, LLMs still face challenges, such as generating inconsistent responses [2], displaying illogical reasoning with out-of-distribution problems [3], and showing overconfidence without understanding their capability limits [4].
Among the many issues, we identify a fundamental category, internal consistency, as central to the core challenges. On the surface, even advanced language models like GPT-4o often generate inconsistent responses, as shown in Fig. 1. At the intermediate level, token selection during decoding, influenced by stochastic sampling methods (Top-k, Top-p, beam search, etc.), can also lead to entirely different answers. At the deepest level,  [5, 6, 7] have shown that specific attention heads in latent layers related to faithfulness exist, meaning different heads may lead to different answers.
Figure 1: GPT-4o provides different answers to the same question. The complete responses can be found in our GitHub repository.
To ensure a model’s internal consistency, several notable approaches have emerged, such as Self-Consistency [2], Self-Refine [8], and Self-Correct [9]. Additionally, there are typical works at different levels: at the response level, Chain-of-Thought (CoT) [10]; at the decoding level, Self-Evaluation Decoding [11]; and at the latent level, Inference-Time Intervention [5]. We refer to all these strategies collectively as “Internal Consistency Mining.”
@@
Internal Consistency Mining refers to developing methods at the response, decoding, or latent level to ensure Large Language Models consistently express their understanding learned from the corpus.
@@ I-A Lack Reasoning and Exhibit Hallucination
Closely related to the internal consistency issue, the challenges of ”lack of reasoning” and ”exhibiting hallucinations” in models also represent persistent concerns. Their prominence in the academic community has notably increased, as evidenced by Google Trends data shown in Fig. 2. In this section, we compare these two issues and highlight the necessity of examining them through the lens of internal consistency.
Figure 2: Relative search interest for the keywords “LLM Hallucination” and “LLM Reasoning” from Google Trends on June 14, 2024.
TABLE I: Relevant Definitions, A Real-World Example, and Benchmark Examples.
@@
Example Query: Which number is mathematically bigger, 1.11 or 1.9? Llama-3.1-8B-Instruct: The number 1.11 is mathematically bigger than 1.9.
@@
As shown in Table I, “reasoning” and “hallucination” have different connotations; however, distinguishing between “lack reasoning” and “exhibit hallucinations” remains challenging. In the real-world example shown in Table I, it is difficult to definitively determine whether “1.11 is greater than 1.9” is due to a hallucination or a lack of reasoning. Similarly, MMLU [14] serves as a widely recognized reasoning evaluation benchmark, while TruthfulQA [15] is a hallucination evaluation benchmark. Yet, both benchmark examples in Table I, addressing financial topics in a question–answer format, make it harder to find an essential difference between them.
Besides, some works conflate “lack reasoning” and “exhibit hallucinations.” For instance, Zhang et al. [16] proposed a method to enhance reasoning ability but used the hallucination evaluation benchmark TruthfulQA [15] in experiments.
Thus, a unified perspective is needed to describe these two closely related phenomena. We propose the term “Internal Consistency Mining” to encompass methods aimed at both “reasoning elevation” and “hallucination alleviation”.
@@ I-B Self-Feedback to Promote Internal Consistency
To enhance a model’s internal consistency, scaling its parameters is the most straightforward approach [17]. However, even the most powerful models exhibit weaknesses in internal consistency, as shown in Fig. 1. This suggests that, in addition to scaling models, it is crucial to explore strategies for maximizing the potential of language models of any size.
So, is there an efficient approach? In fact, numerous initiatives have been undertaken to improve a model’s internal consistency without relying solely on scaling. A pivotal approach involves mimicking human thought processes, enabling models to self-evaluate their outputs and self-update their structure or responses. Notable examples include Self-Consistency [2], which prompts the model to generate multiple answers to check for consistency (Self-Evaluation), and then use a majority voting strategy to select the final answer (Self-Update), thereby enhancing reasoning capabilities. Another example is Self-Contradict [18], which induces models to generate diverse content and checks for contradictions (Self-Evaluation), allowing the model to resolve contradictions autonomously (Self-Update) to reduce hallucinations.
Moreover, during Self-Evaluation, it is possible to not only inspect the model’s responses but also examine its logits and the latent states. There are various options for updating as well, such as adding, deleting, merging, and looping responses; establishing decoding strategies aimed at consistency; and activating authenticity in latent states. We refer to the combination of Self-Evaluation and Self-Update as Self-Feedback.
I-C Related Surveys
Surveys [19, 20, 21] are similar to ours. We present a straightforward comparison in Table II.
@@ I-C Related Surveys
A Survey on Self-Evolution of Large Language Models [19] covers literature on LLMs generating their own training data and using multi-agent approaches for iterative optimization. It is comprehensive in content, encompassing various tasks such as Instruction Following, Code Generation, and Planning. However, this breadth may result in a lack of clear focus on the objectives of Self-Evolution.
Automatically Correcting Large Language Models: Surveying the Landscape of Diverse Automated Correction Strategies [20] focuses on Self-Correction, where models correct their own errors. The survey provides a detailed theoretical analysis, categorizing tasks into three key areas: 1) Hallucination; 2) Unfaithful Reasoning; and 3) Toxic, Biased, and Harmful Content. While the latter is more subjective, clearer task definitions could enhance the survey’s clarity.
When Can LLMs Actually Correct Their Own Mistakes? A Critical Survey of Self-Correction of LLMs [21] questions whether models can truly Self-Correct, focusing on cases where feedback is textual and partially external. This narrow scope limits the comprehensiveness of the survey’s conclusions, which we further analyze in Section IX.
@@
    Internal consistency perspective. We offer an in-depth review of LLMs’ internal consistency, examining its phenomena, formalization, status quo, etc. Furthermore, we introduce the task of Internal Consistency Mining, providing a unified perspective for reasoning elevation and hallucination alleviation tasks.
@@
    Self-Feedback theoretical framework. Our framework includes Self-Evaluation, Consistency Signal Acquisition, and Self-Update. Characterized by its simplicity and comprehensiveness, this framework is poised to inspire further research. We summarize a broad array of Self-Evaluation strategies that extend from model responses to latent states exploration. These strategies allow us to capture a diverse range of Feedback Signals, extending beyond the scalar, textual, and external signals discussed in other surveys, to include contrastive signals.
@@
I-D Structure of the Survey
As shown in Fig. 3, our research begins with the existing problem of low internal consistency in LLMs (Section II-C). Specific manifestations of low internal consistency include poor reasoning capabilities in question-answering (QA) scenarios and hallucinations in free-form generation (Section I-A). From a causal perspective, elements contributing to low internal consistency include inadequate latent reasoning, the snowball effect of hallucinations, and the stochastic parrot hypothesis (Section II-D). We formalize internal consistency as the sampling-based consistency of model expressions across different layers (Section II-A). This involves enhancing response, decoding, and latent consistency (Sections II-AII-B).
To improve internal consistency, we propose Internal Consistency Mining across these layers. While scaling up the model is an intuitive solution, it comes with various cost-related challenges (Section I-B). Thus, we focus on the Self-Feedback theoretical framework, which mainly includes Self-Evaluation, Consistency Signal Acquisition, and Self-Update. Models obtain different forms of internal consistency signals through Self-Evaluation, and subsequently use these signals to Self-Update either responses or the model itself (Section III). We explore six lines of work in Consistency Signal Acquisition (Section IV) and seven lines of work utilizing the Self-Feedback framework, divided into three lines dedicated to reasoning elevation (Section V) and four lines aimed at hallucination alleviation (Section VI).
@@ I-E Out-of-scope Topics
  • Papers reviewed in this work mainly employ the Self-Feedback framework and show improvements in the internal consistency. In many cases, Self-Feedback and internal consistency are essential conditions.
  • This survey focuses exclusively on internal consistency and does not explore the interaction between internal and external consistencies. Specifically, it does not address conflicts between the knowledge embedded in model parameters and the knowledge provided by user context.
  • In line with many related surveys, our focus is on the model’s self-awareness, self-assessment, self-correction, etc. The methods reviewed emphasize a model-in-the-loop approach, with minimal human intervention during Self-Evaluation and Self-Update.
  • While retrieval-augmented generation (RAG) is recognized for mitigating external hallucinations [22], this paper does not actively discuss RAG. Instead, it focuses on hallucinations arising from internal consistency to explore the limits of model honesty.
@@ II-A Formulation
Consistency is a critical term in logic, referring to a system where no two statements contradict each other [23]. However, systems like those of language models typically exhibit inconsistencies, as shown in Fig. 1. To better define the internal consistency, we utilize a sampling-based approach to model expressions in LLMs [24]. In addition, Table III provides explanations of some notations frequently used in this paper.
@@
    The rewritten query can lead to significant changes in the answer [25].
@@
  • Expression from Latent Layer (tensor). Expression at this layer encompasses the different activation of attention heads and latent states across the model’s architecture, contributing to diverse outputs.
For the expression type $e$, the expression distribution produced by $\mathcal{M}$ in response to $\bm{x}$ can be defined as follows:
@@
By sampling from this distribution, we can obtain a sampling set with potentially repeated elements:
@@
We will comprehensively discuss existing methods for acquiring consistency signals in Section IV. Those methods may be more helpful.
@@ II-B The Hourglass Evolution of Internal Consistency
Response Layer. We used Top-p sampling with a fixed temperature to sample five times. To induce diverse responses, CoT prompting was enabled. We observed the model’s final textual choices during free generation. One example output is: “Let’s think step by step. There is one period at the end of the first part, … So, there are 3 periods in total.” The resulting sampling set is $\mathcal{Y}_{\text{response}}=\{5,3,3,3,3\}$.
@@
The experimental results are also shown in Fig. 5. We observed that the model’s answer consistency follows an “hourglass evolution” pattern, starting from the lower to higher layers at the latent level, passing through the intermediate decoding level, and finally reaching the response level.
@@
We analyze this phenomenon as follows. In the latent state, since the forward propagation is not yet complete, the attention heads near the bottom layers may tend to choose answers randomly. In contrast, the attention heads near the top layers can continually accumulate knowledge due to residual connections, leading to a gradual convergence in judgment. During the decoding phase, all decoding strategies tend to select the token with the higher probability, thus maintaining high certainty. However, at the response stage, greater variability appears. When the LLM generates the first token, it has already conducted reasoning (namely, the latent reasoning [26]) and made an initial judgment of the answer. However, during the response phase, the output tokens such as “I’m willing to help.” can interfere with the model’s initial reasoning and preliminary judgment, leading to a collapse of latent reasoning.